In the Bohr model of the hydrogen atom, the electron in the n = 11 level moves in a circular orbit of radius 6.40 x 10-9 m arHello, I tried 3.1e13 and 3.1e14

but they says it wrong!
Answer #1

c] angular frequency w = v/r = 1.99e5/6.4e-9

= 3.11e13 rad/s

By other method, angular momentum = I*w

11h/2pi = mr^2*w

w = [11h/2pi]/(mr^2)

= [11*6.626e-34/(2pi)]/[9.1e-31*6.4e-9^2]

= 3.11e13 rad/s answer